OpenStudy (anonymous):

You drive a car that holds 15 gallons of gasoline. Your car uses 0.03 gal per mi. Which equation best represents this information?

OpenStudy (tcp2315):

y = 15 - 0.03x where x is the miles driven and y is the amount of gas

OpenStudy (anonymous):

y = 15- .03 x

OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped two degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x.